Using the data from 15 automobile accidents, the correlation coefficient between the combined speeds of the cars (x) in an accident and the amount of damage done (y) is 0.7831. The regression equation for the two variables is y = 801.518 + 162.845x. 6-7.2: Is this a significant correlation? 6-7.3: If the answer to part (a) is YES, then predict the amount of damage done in an accident in which the combined speeds of the car involved was 100 mph. 6-7.4: A car traveling at 28 mph is involved in an accident with another car which did $9,780 in damages to the cars. Estimate how fast the other car was driving
Answer #1

6-7.2: Is this a significant correlation?

Yes

6-7.3: If the answer to part (a) is YES, then predict the amount of damage done in an accident in which the combined speeds of the car involved was 100 mph.

y = 801.518 + 162.845x

Put x = 100

y = 801.518 + 162.845*100

y = 17086.018

6-7.4: A car traveling at 28 mph is involved in an accident with another car which did $9,780 in damages to the cars. Estimate how fast the other car was driving

y = 801.518 + 162.845x

Put y = 9780

9780 = 801.518 + 162.845x

x = 55.13514078

The other car speed = 55.13514078 - 28 = 27.13514078

Round off the decimals as per the requirement.

Please give me a thumbs-up if this helps you out. Thank you!